OpenStudy (anonymous):

Which is equivalent to (x2 - 4)/(x2 - 4x +4)?

OpenStudy (luigi0210):

\[\frac{ x^2-4 }{ x^-4x+4 }=\frac{ (x+2)(x-2) }{ (x-2)(x-2) }=\frac{ x+2 }{ x-2 }\]

OpenStudy (anonymous):

Ohhhh, that's how you do it! I totally forgot. Thankyou so much.

OpenStudy (luigi0210):

Yup :P